# The Scene Language: Representing Scenes with Programs, Words, and Embeddings

### v1

@@ 2.3 Representing Scenes with Generative Model Latents
Special Forms
(bind <word > <entity-func>)// Defines and binds an entity function
Table 2: The Domain-Specific Language. The following table contains the DSL specification used to define our representation. Built-in data types (e.g., Float), functions (e.g., car and cdr), special forms (lambda), and conditionals (if) are omitted; <START> denotes program starts, ::= denotes rewriting rules; :: denotes type annotation; _:: denotes type annotation for an anonymous formal parameter, * denotes one or more expressions of the same type. The full DSL and explanations of the macros are deferred to Appendix B.
3 The Scene Language
We aim to design a visual scene representation that encodes the structure, semantics, and visual content of scenes. Towards this goal, we propose the Scene Language, which represents a scene with three components: a program that encodes scene structure by specifying the existence and relations of scene components, which we will refer to as entities; words in natural language that denote the semantic group of each entity in the scene; and neural embeddings that pertain the low-level visual details and identities of the entities by permitting an expressive input parameter space. In the following, we will first give a formal definition of the representation (Section 3.1), and then introduce a domain-specific language (DSL) (Section 3.2) as its realization.
3.1 Formal Definition
$\Phi(s):=\left(W,P,Z\right).$ (1)
In the following, we will expand on how functions from $P$ are defined, followed by the program execution procedure to compute the represented scene $s$. Notations are summarized in Table 1.
Entity Function Definitions.
$\text{where}\;\;h^{(i)}$ $=f_{w^{(i)}}(z^{(i)},\gamma^{(i)})\;\;\text{is a sub-entity of $h$ and}\;\;t^{(i)}(z)\;\;\text{specifies its pose.}$.
Program Execution.
3.2 The Scene Language as a Programming Language
We now concretize the definition in Section 3.1 with a domain-specific language (DSL) specified in Table 2. To define entity functions in the DSL, we introduce macro operations union for $\Psi_{\text{union}}$, union-loop which calls union on entities evaluated in a for-loop, and transform for $\Psi_{\text{transform}}$. We use these four macro operations and function calls of dependent functions to define entity functions. Entity functions are identified with the associated words in the DSL via two special forms: bind, which binds an entity function $f_{w}$ to word $w$, and retrieve, which retrieves $f_{w}$ given $w$ and applies $f_{w}$ on actual embedding parameters. If $w$ is never bound to a function, it corresponds to an entity function with no sub-entities (i.e., $N=0$ in Eq. 2), in which case call returns a primitive entity with no children.
The data type of an entity $h$ (Eq. 2) is denoted as Entity. It stores two data fields Word and Embedding, describing the entity’s semantic group and identity respectively, and stores each children entity together with its pose in the frame of $h$. In particular, Embedding captures the visual details and requires a highly expressive representation, e.g., neural embeddings. We employ the textual embedding space of OpenCLIP-ViT/H (Ilharco et al., 2021) for parameterization, denoted as $\mathcal{Z}_{\text{CLIP}}$. It offers the advantage that embeddings can be either encoded directly from natural language or inferred from images with Textual Inversion (Gal et al., 2023). Table 1 summarizes the expressions and data types in accordance with the notations from Section 3.1.
Renderer Examples
Primitive-based renderer Light transport simulation Shape and BRDF parameters LM inference
Asset-based renderer Ray tracing Asset metadata LM inference
SDS-based renderer Gaussian splatting 3D Gaussian parameters SDS optimization
T2I model Model feed-forward pass Text embeddings in $\mathcal{Z}$ CLIP text encoding
Table 3: Graphics Renderers Examples.
4 Rendering
Renderer Specifications.
Renderer Instantiations.
An example renderer instantiation is with Score Distillation Sampling (SDS) (Poole et al., 2023) guidance, where $\Theta$ is a differentiable 3D representation, and we specify $g_{\text{reparam}}$ as follows. Recall that an entity is associated with a Word value, e.g., moai for one statue from Figs. 1 and 3, and an Embedding value, e.g., <z2>. For each primitive entity (i.e., an entity with no children), given these two value fields together with the fields of its ancestors, we use a manually specified language template $c$, or a <z2> moai, in the style of <z1>, 3D model in this example, to embed them into $z=g_{\text{CLIP}}(c)\in\mathcal{Z}_{\text{CLIP}}$ where $g_{\text{CLIP}}$ is the pre-trained CLIP text encoder. Then, $g_{\text{reparam}}:\mathcal{Z}_{\text{CLIP}}\rightarrow\Theta$ corresponds to the SDS-guided optimization to find a solution in $\Theta$ that aligns with the input condition $z\in\mathcal{Z}_{\text{CLIP}}$. Output 3D scenes can be personalized by editing embeddings, e.g., <z1> which controls the global style in Fig. 1.
For the underlying 3D representation, we use 3D Gaussian Splatting (Kerbl et al., 2023) where images are rendered by splatting a set of 3D Gaussians onto the image plane; other differentiable 3D representations such as neural fields will also be suitable. We base our implementation on GALA3D (Zhou et al., 2024c), and use MVDream (Shi et al., 2024) and a depth-conditioned ControlNet (Zhang et al., 2023b) for guidance.
We refer to the renderer above as the Gaussians renderer. Other possible renderers include primitive-based renderers, such as Mitsuba (Jakob et al., 2022) with graphics primitives of cubes, spheres, and cylinders; asset-based game engines, such as MineCraft (Mojang Studios, 2009); and feed-forward inference of layout-conditioned text-to-image (T2I) diffusion models, such as MIGC (Zhou et al., 2024a), which achieves 2D bounding box conditioning by controlling attention layers from Stable Diffusion (Rombach et al., 2022)). Table 3 shows a summary; details are in Appendix C.
Figure 4: Text-Conditioned Scene Generation. The bottom shows input texts. Ours produce more accurate outputs than using no intermediate representation (MVDream) or scene graph (GraphDreamer).
Figure 5: Scene Editing with Language Instructions. The program structure from our representation is highly interpretable, which benefits user editing. The bottom of each row shows initial scene descriptions and editing instructions in the format of user text prompts. We prompt an LM to infer the initial Scene Language as well as the edits (shown with texts in red), and show image renderings with two renderers.
5 Inference via Pre-trained Language Models
LMs have shown remarkable capability in code generation with common programming languages such as Python. In our implementation, we prompt LMs to generate Python scripts. We prompt the LM with the input condition, i.e., a scene description in texts or an image; a Python script of helper functions converted from the DSL in Table 2; and an example script using the helper functions. We use Claude 3.5 Sonnet (Anthropic, 2024) for all experiments for our method and LM-dependent baselines. Full language prompts are listed in Appendix D.
Function arguments in the LM-generated scripts, which are numeric values or string tokens, are converted to embeddings from $\mathcal{Z}_{\text{CLIP}}$ (Section 3.2) using language templates and the CLIP text encoder $g_{\text{CLIP}}$. For example, in the raw LM output, function calls for white pieces in Fig. 2 have input attribute {"color":(.9,.9,.9)}, and we prompt LM to describe the color value as a word, and feed the word into $g_{\text{CLIP}}$ to compute <z68>. For image-conditioned tasks, for each primitive entity in the execution output of $P$, we first use GroundingSAM (Kirillov et al., 2023; Ren et al., 2024) to segment out the region defined by the word associated with the entity. We then use Textual Inversion (Gal et al., 2023) to optimize an embedding to reconstruct the cropped image with the diffusion model training objective. The full process is deferred to Section E.1.
6 Applications
We apply the method from Section 5 to the tasks of text-conditioned 3D scene generation (Section 6.1) and editing (Section 6.2), image-conditioned scene generation (Section 6.3), and 4D scene generation (Section 6.4).
6.1 Text-Conditioned Scene Generation
Baselines.
To evaluate the proposed representation, we compare our inference pipeline with 3D scene generation methods using alternative intermediate representations, e.g., scene graph. In particular, we compare with GraphDreamer (Gao et al., 2024) as an exemplar approach, which generates scene graphs from input texts via LM prompting and then synthesizes scenes conditioned on the graphs via SDS guidance. We further ablate the role of structural representation in this task by comparing ours with the backbone of our SDS-based renderer, MVDream (Shi et al., 2024), as a direct scene generation approach. Full implementation details in Section E.2.
Results.
Text-conditioned scene generation results rendered with the SDS-based renderer are shown in Fig. 4. Compared to the direct 3D scene generation method MVDream, our approach is compositional and adheres more closely to input prompts in scenes involving multiple objects. Compared to a scene graph representation, where entity relations are restricted to be between two objects and are bottlenecked by the coarseness of natural language descriptions, e.g., “aligned in a row”, a program-based representation offers more flexible and precise specifications for relations, e.g., the particular coke can arrangement in Fig. 4. This brings the practical benefit of offloading the burden of generating scenes involving complex entity relations from the T2I model (used for SDS guidance in both ours and GraphDreamer) towards LM, leading to accurate and detailed generation results.
Methods Alignment Counting
GraphDreamer $3.56$±$\pm 7.38$ 0.11
MVDream $10.79$±$\pm 12.83$ 0.11
Ours 85.65±$\pm\textbf{13.71}$ 1.0
Table 4: Quantitative Evaluation. We perform a user study to compare with prior methods on the text-conditioned 3D generation task and report the percentages of user preferences for prompt alignment. We also report the counting accuracy (00 for inaccurate and $1$ for accurate). Results are averaged across $9$ scene prompts and $103$ users; $\pm$ denotes standard deviation.
To quantitatively compare our method with baselines, we conduct a user study (Section E.3) where users are asked to choose one of the three animated scenes, synthesized by ours and two baselines in a randomized order, that aligns the best with the text prompt for the scene. We further report whether the synthesized scenes have the correct object count. As shown in Table 4, our method achieves a more favorable prompt alignment than the baselines and has a clear advantage in counting accuracy.
6.2 Text-Instructed Scene Editing
Scenes synthesized from our proposed representation can further be edited by prompting LM with its previously generated script and an editing instruction in natural language instruction. Results are shown in Fig. 5. Our representation provides an interpretable and intuitive interface for scene editing, as functions have explicit semantic meanings associated with words, and function reuse significantly improves program readability. Furthermore, since the structure of programs reflects the structure of scenes, editing program parameters leads to changes in the scenes while preserving the original structure, e.g., the circular arrangement of staircases in Fig. 5. Desirable editing effects involving multiple primitives, or all staircases in this example, can be effectively achieved via only minor changes in the program space. Finally, the program structure itself, such as the function header in the Jenga set example, can be adjusted for editing to achieve localized edits that only affect relevant parts of the scene.
Figure 6: Image-Conditioned Scene Generation. Both ours and GraphDreamer parse an input image to semantic entities. Programs from ours encode additional scene structure, e.g., repetitions, and specify geometric relations among entities more precisely. Embeddings from ours further enable visual identity preservation in the renderings.
The compositionality of our representation directly benefits localized scene editing. In comparison, MVDream from Section 6.1 does not apply to this task, as the entire scene is parameterized with a single 3D representation. Precisely encoding the geometric relations of scene components further enhances the controllability of generated scenes. In comparison, GraphDreamer represents the binary relation of scene components with coarse language descriptions; therefore, it does not apply to editing tasks involving precise geometric controls, as shown in the first example from Fig. 5.
6.3 Image-Conditioned Scene Generation
The proposed representation can be used for image parsing and generating 3D scenes consistent with the parsed image structure and content. We compare our representation with scene graphs by comparing our method with GraphDreamer as shown in Fig. 6. Our representation explicitly encodes semantic components parsed from input images, high-level scene structures, such as the repetition of Coke cans, and visual details, such as the glass bottles with particular shapes and colors. Our method preserves both structure and visual content from input images. In contrast, GraphDreamer only reconstructs semantics from input images; it leaves out entity poses and identities due to the information loss in the intermediate scene graph representation.
6.4 Text-Conditioned 4D Scene Generation
We apply the inference method from Section 5 to generate 4D scenes. The 4D scene representation in this task is identical to the definition in Eq. 1, except that there is an additional 4D entity function in the program $P$. The corresponding DSL extends from Table 2 as specified in Appendix B.
Allowing for a flexible set of primitive entities is crucial to make our representation suitable for generating 4D scenes of different scales, including objects with moving parts (e.g., the wind turbine from Fig. 8) and scenes with moving objects (e.g., the carousel). Specifically, primitives have granularity adapted to the particular scene being represented, instead of being chosen from a fixed set (Tian et al., 2019; Sharma et al., 2018) or object-centric as in scene graphs (Johnson et al., 2015).
Figure 7: Text-Conditioned 4D Scene Generation. The proposed representation may also capture scene dynamics and be applied for synthesizing 4D scenes. It explicitly represents the temporal correspondence of an entity in a dynamic scene. Each colored trajectory denotes tracking of a temporally moving point. See our project page for animations.
Figure 8: Visualizations of Discriminative Maps. The proposed representation contains semantics information for scene components, visualized using semantic segmentation shown in (b). It is compositional and directly informs instance segmentation (c). Furthermore, it specifies the dense correspondence across repeated entities (d).
Moreover, the hierarchical scene structure encapsulated by our program-based representation makes it possible to represent 4D scenes compactly, serving as a regularization for generation output. Multiple entities, such as outputs from the function horse from the carousel scene in Fig. 8, can be grouped into one function horses and thereby share the same temporal transformation. Writing composible functions for entity grouping effectively reduces the dimension of the temporal motion space and improves motion fidelity. See the project page for better visualizations.
6.5 Different Graphics Renderers
Figure 9: Renderings Across Graphics Renderers. Different renderers produce renderings that adhere to the same representation and are therefore visually aligned, while each exhibits a different imaging style. Text inputs are shown at the bottom of the subfigures.
The same program can be rendered with different renderers described in Section 4, showing the versatility of the proposed representation. The results are shown in Fig. 9 with the same experiment setup as in Section 6.1.
6.6 Visualization of Discriminative Information
As shown in Fig. 8, several pieces of discriminative information can be directly obtained with the proposed Scene Language: semantic maps in Fig. 8b, as words represent per-entity semantics; instance segmentation in Fig. 8c, as the representation is compositional with separable instances; correspondence of the repeated instances in Fig. 8d, as programs specify repetitions existing in a scene; dense temporal correspondence for 4D scenes, as shown in Fig. 8.
7 Conclusion
We have introduced a visual scene representation, termed the Scene Language, which encodes three key aspects of visual scenes: scene structure, such as hierarchy and repetition, specified via programs; semantics of individual components succinctly summarized via words; and identities of each component precisely captured via neural embeddings. We formalize the representation as a programming language defined using a DSL. We show that the Scene Language can be efficiently inferred from both text and image inputs using pre-trained language models. Once the program is executed, the resulting scene can be rendered into images using a variety of graphics renderers. Compared with existing methods, our Scene Language produces 3D and 4D scenes with significantly higher fidelity, preserves complex scene structures, and enables easy and precise editing.
Appendix A Overview
The supplementary contains the details for the representation definition (Appendix B), graphics renderers (Appendix C), and experiments (Appendix E). Please refer to main text to see how they are integrated.
Appendix B Domain-Specific Language
The complete DSL includes the ones listed in Table 2, with additional macros for computing transformation matrices as listed in Table 5, and grammar for 4D scenes as listed in Table 6.
We explain the four macros form Table 2 as follows. Their type annotations are listed in Table 5.
  • Macro call retrieves the <entity-func> bound to the input word, applies the function on the input embeddings, and outputs a data object of type Entity. Specifically, (car embedding-list) is the embedding of the output entity corresponding to $z$ from Eq. 2, and (cdr embedding-list) is the list of embeddings of its descendent entities corresponding to $\gamma$.
  • Macro union-loop applies union in a for loop.
  • Macro transform pairs an entity with its pose.
Data Types
Vector ::= Array[Float]// Vector in $\mathbb{R}^{3}$
Macros
@@ Appendix B Domain-Specific Language
Table 5: The Domain-Specific Language includes the definitions from Table 2 and the transformation-related macros from this table.
Grammar
4D-entity-func ::= (lambda () create-entity-list)// Define a function that outputs a 4D scene
create-entity-list ::= (list create-entity*)// Represent a 4D scene as a temporal list of entities
Table 6: The Domain-Specific Language for 4D scenes. * indicates one or more expressions.
Appendix C Details of Graphics Renderers
This section expands the instantiation of three graphics renderers from Section 4 in detail. For each rendere, we will discuss its parameter space $\Theta$ and $\mathcal{T}$, renderer $\mathcal{R}$, and the reparameterization function $g_{\text{reparam}}$.
C.1 SDS-Based Renderer
@@
Following Kerbl et al. (2023), a single Gaussian is defined by
@@
We assume that diagonal entries of the scaling matrix $S_{t}$ are all positive, and therefore $t$ is invertible. Combining the above gives
@@
The rendering operation $\mathcal{R}$ to convert an entity consisting of $N$ primitives, $\{(\theta_{j},t_{j})\}_{j=1}^{N}=\{(K_{j},\{\phi_{i}^{j}\}_{i=1}^{K_{j}},t_{j})\}_{j=1}^{N}$, to the image space simply amounts to rendering all post-transformation 3D Gaussians in the scene, $\{\tilde{t}_{j}(\theta_{j})\}_{j}:=\{\tilde{t}_{j}(\phi_{i})\}_{i,j}$, following the projection and blending process from Kerbl et al. (2023).
@@
Recall that $g_{\text{reparam}}$ aims to obtain 3D Gaussian primitive parameters for per-primitive conditional embeddings $\{z_{j}\}_{j=1}^{N}$ and global condition $z_{\text{global}}$, where $z_{j}=g_{\text{CLIP}}(c_{j})$ is explained in Section 4, and $z_{\text{global}}=g_{\text{CLIP}}(c_{\text{global}})$ is computed from a global scene description in texts, $c_{\text{global}}$. We now expand Section 4 to describe the optimization process of $g_{\text{reparam}}$ in detail.
We write the SDS objective originally proposed in Poole et al. (2023) as follows:
$g(\psi;z,\hat{\epsilon}):=\nabla_{\psi}\mathcal{L}_{\text{SDS}}(x=\mathcal{R}(\psi);z,\hat{\epsilon})=\mathbb{E}_{\eta\sim\mathcal{U}(0,1),\epsilon\sim\mathcal{N}(0,I)}\left[w(\eta)(\hat{\epsilon}(\alpha_{\eta}x+\alpha_{\eta}\epsilon,z,\eta)-\epsilon)\frac{\partial x}{\partial\psi}\right],$
@@
With the notations from above, for entity $\{(\theta_{j},t_{j})\}_{j=1}^{N}$, let
($\mathcal{L}(\{z_{j}\}_{j},z_{\text{global}},\{t_{\text{init},j}\}_{j}):=$ $\mathcal{L}_{\text{SDS}}(\{\tilde{t}_{j}(\theta_{j})\}_{j};z_{\text{global}},\hat{\epsilon}_{\text{ControlNet}})+\sum_{j}\mathcal{L}_{\text{SDS}}(\theta_{j};z_{j},\hat{\epsilon}_{\text{MVDream}})$
@@
where $\mathcal{L}_{\text{reg}},\mathcal{L}_{\text{layout}}$ are regularization terms following the definition from Zhou et al. (2024c) and StopGrad stops gradients from backpropagation. Here, $\mathcal{L}_{\text{reg}}$ penalizes Gaussian ellipsoids that are too long, and $\mathcal{L}_{\text{layout}}$ penalizes Gaussians that lie outside the intial bounding box specified by $t_{\text{init}}$.
@@
$g_{\text{reparam}}=\arg\min_{\{(\theta_{j},t_{j})\}_{j=1}^{N}}\mathcal{L}.$
@@ C.2 Mitsuba Renderer
For this renderer, $\Theta$ is the parameter space for three types of graphics primitives supported by Mitsuba: cube, sphere, and cylinder, as specified in the function header for primitive_call in Section D.1. $\mathcal{T}$ is the 3D affine transformation space.
@@
We use the path tracer with maximum depth 8 implemented in Mitsuba.
@@
Since we directly prompt LM to generate Mitsuba primitive parameters in its outputs as specified in Section D.1, the function parameters from raw LM outputs are already in the parameter space $\Theta$ and are directly used for rendering, instead of being encoded into CLIP embeddings $z\in\mathcal{Z}_{\text{CLIP}}$.
C.3 Minecraft Renderer

### v2

@@ B.1 Domain-Specific Language
<4D-entity-func> ::= (lambda () <create-entity-list>)// Define a function that outputs a 4D scene
<create-entity-list> ::= (list <entity>*)// Represent a 4D scene as a temporal list of entities
@@
Special Forms
(bind <word > <entity-func>)// Defines and binds an entity function
Table 5: The Domain-Specific Language. The table contains the DSL specification used to define our representation. Built-in data types (e.g., Float), functions (e.g., car and cdr), special forms (lambda), and conditionals (if) are omitted; <START> denotes program starts, ::= denotes rewriting rules; :: denotes type annotation; _:: denotes type annotation for an anonymous formal parameter, * denotes one or more expressions of the same type.
B.2 Details of Graphics Renderers
This section expands the instantiation of three graphics renderers from Sec. 4 in detail. For each renderer, we will discuss its parameter space $\Theta$ and $\mathcal{T}$, rendering operation $\mathcal{R}$, and the reparameterization function $g_{\text{reparam}}$.
B.2.1 SDS-Based Renderer
@@
Following [18], a single Gaussian is defined by
@@
We assume that diagonal entries of the scaling matrix $S_{t}$ are all positive, and therefore $t$ is invertible. Solving for the above equation gives
@@
The rendering operation $\mathcal{R}$ to convert an entity consisting of $N$ primitives, $\{(\theta_{j},t_{j})\}_{j=1}^{N}=\{(K_{j},\{\phi_{i}^{j}\}_{i=1}^{K_{j}},t_{j})\}_{j=1}^{N}$, to the image space simply amounts to rendering all post-transformation 3D Gaussians in the scene, $\{\tilde{t}_{j}(\theta_{j})\}_{j}:=\{\tilde{t}_{j}(\phi_{i})\}_{i,j}$, following the projection and blending process from [18].
@@
Recall that $g_{\text{reparam}}$ aims to obtain 3D Gaussian primitive parameters for per-primitive conditional embeddings $\{z_{j}\}_{j=1}^{N}$ and global condition $z_{\text{global}}$, where $z_{j}=g_{\text{CLIP}}(c_{j})$ is explained in Sec. 4, and $z_{\text{global}}=g_{\text{CLIP}}(c_{\text{global}})$ is computed from a global scene description in texts, $c_{\text{global}}$. We now expand Sec. 4 to describe the optimization process of $g_{\text{reparam}}$ in detail.
We write the SDS objective originally proposed in [25] as follows:
$\nabla_{\psi}\mathcal{L}_{\text{SDS}}(x=\mathcal{R}(\psi);z,\hat{\epsilon})$
$=$ $\mathbb{E}_{\eta\sim\mathcal{U}(0,1),\epsilon\sim\mathcal{N}(0,I)}\left[w(\eta)(\hat{\epsilon}(\alpha_{\eta}x+\alpha_{\eta}\epsilon,z,\eta)-\epsilon)\frac{\partial x}{\partial\psi}\right],$
@@
For entity $\{(\theta_{j},t_{j})\}_{j=1}^{N}$, let
($\mathcal{L}(\{z_{j}\}_{j},z_{\text{global}},\{t_{\text{init},j}\}_{j})$
::=:= $\mathcal{L}_{\text{SDS}}(\mathcal{R}(\{\tilde{t}_{j}(\theta_{j})\}_{j});z_{\text{global}},\hat{\epsilon}_{\text{ControlNet}})$
$+\sum_{j}\mathcal{L}_{\text{SDS}}(\mathcal{R}(\theta_{j});z_{j},\hat{\epsilon}_{\text{MVDream}})$
@@
where $\mathcal{L}_{\text{reg}},\mathcal{L}_{\text{layout}}$ are regularization terms following the definition from [44] and StopGrad stops gradients from backpropagation. Here, $\mathcal{L}_{\text{reg}}$ penalizes Gaussian ellipsoids that are too long, and $\mathcal{L}_{\text{layout}}$ penalizes Gaussians that lie outside the intial bounding box specified by $t_{\text{init}}$.
@@
$g_{\text{reparam}}=\operatorname*{arg\,min}_{\{(\theta_{j},t_{j})\}_{j=1}^{N}}\mathcal{L}.$
@@ B.2.2 Mitsuba Renderer
For this renderer, $\Theta$ is the parameter space for three types of graphics primitives supported by Mitsuba: cube, sphere, and cylinder, as specified in the function header for primitive_call in Sec. F.1.1. $\mathcal{T}$ is the 3D affine transformation space.
@@
We use the path tracer with maximum depth 8 implemented in Mitsuba. In this work, we use Mitsuba as a generic graphics engine and do not take advantage of its differentiability.
@@
Since we directly prompt LM to generate Mitsuba primitive parameters in its outputs as specified in Sec. F.1.1, the function parameters from raw LM outputs are already in the parameter space $\Theta$ and are directly used for rendering, instead of being encoded into CLIP embeddings $z\in\mathcal{Z}_{\text{CLIP}}$.
B.2.3 Minecraft Renderer